# FinTagging: An LLM-ready Benchmark for Extracting and Structuring Financial Information

### v1

@@
Abstract
We introduce FinTagging, the first full-scope, table-aware XBRL benchmark designed to evaluate the structured information extraction and semantic alignment capabilities of large language models (LLMs) in the context of XBRL-based financial reporting. Unlike prior benchmarks that oversimplify XBRL tagging as flat multi-class classification and focus solely on narrative text, FinTagging decomposes the XBRL tagging problem into two subtasks: FinNI for financial entity extraction and FinCL for taxonomy-driven concept alignment. It requires models to jointly extract facts and align them with the full 10k+ US-GAAP taxonomy across both unstructured text and structured tables, enabling realistic, fine-grained evaluation. We assess a diverse set of LLMs under zero-shot settings, systematically analyzing their performance on both subtasks and overall tagging accuracy. Our results reveal that, while LLMs demonstrate strong generalization in information extraction, they struggle with fine-grained concept alignment, particularly in disambiguating closely related taxonomy entries. These findings highlight the limitations of existing LLMs in fully automating XBRL tagging and underscore the need for improved semantic reasoning and schema-aware modeling to meet the demands of accurate financial disclosure. Code is available at our GitHub repository [^1] https://github.com/The-FinAI/FinTagging and data is at our Hugging Face repository [^2] https://huggingface.co/collections/TheFinAI/fintagging-68270132372c6608ac069bef.
1 Introduction
@@
Automated tagging is essential in financial reporting, converting structured content such as tables and text into machine-readable data by linking each number to its meaning. Globally, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, which provide essential information for investors, regulators, and analysts. However, inconsistent terminology poses challenges for reliable interpretation. To address this, the eXtensible Business Reporting Language (XBRL) was introduced in 1999 as a global standard for machine-readable financial data richards2006introduction . By applying structured tags, XBRL transforms fragmented disclosures into a consistent format, enabling seamless data exchange, cross-entity comparisons, and large-scale analysis. Yet, as shown in Figure 1, accurately tagging over 2,000 numerical facts per report to more than 10,000 standardized concepts remains labor-intensive and error-prone. In 2023 alone, over 6,500 tagging errors were identified across 33,000 SEC filings [^3] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
Figure 1: An example of Financial Tagging.
@@
Existing approaches to automated XBRL tagging rely on task-specific models and pretrained language models (PLMs) loukas2022finer ; sharma2023financial , but they require costly annotations, struggle to generalize to new filings, and perform poorly as taxonomies scale. In contrast, recent large language models (LLMs) show strong zero- and few-shot reasoning without task-specific supervision wei2022chain ; cao2025development ; wang2025ordrankbennovelrankingbenchmark ; peng2025plutusbenchmarkinglargelanguage , yet their application to XBRL tagging remains underexplored due to two major limitations (Table 1). First, task framing is oversimplified: prior benchmarks like FiNER loukas2022finer and FNXL sharma2023financial treat tagging as extreme multi-class classification over flat label sets, offering little context for disambiguating fine-grained financial facts. These datasets typically cover only 1k+ concepts, leaving most of the 10k+ US-GAAP taxonomy untested. As shown in Table 8, SOTA LLMs such as DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt achieve 0.0 precision, recall, and F1 under this setting. Second, structured data is ignored: existing datasets exclude tables, despite their central role in financial reporting. Many key facts appear in structured formats, as shown in recent financial QA benchmarks chen2021finqa ; zhu2021tat ; mathew2021docvqa , making this omission a key gap between benchmarks and real-world tagging needs.
@@
Table 1: Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities. “QA” means the question-answering task. “Num. Reasoning” indicates the numerical reasoning. “Struct. IE” denotes the structured information extraction.
Benchmark Scenario Data Source Task Modality #Entity Label #Taxonomy Label Num. Reasoning? Struct. IE? Concept Linking?
FinQA chen2021finqa Decision making SEC 10-K QA text/table 0 0
ConvFinQA chen2022convfinqa Decision making SEC 10-K QA text/table 0 0
TAT-QA zhu2021tat Financial analysis Chinese financial reports QA text/table 0 0
@@
DocVQA mathew2021docvqa Enterprise automation Financial document images QA text/image 0 0
@@
FiNER-ORD shah2023finer Financial tagging Chinese financial disclosures Classification text 49 0
@@
FinRED sharma2022finred Knowledge graph construction English financial news Classification text 38 0
FiNER loukas2022finer XBRL tagging SEC 10-K Extreme classification text 0 139
FNXL sharma2023financial XBRL tagging SEC 10-K Extreme classification text 0 2,800
@@
FinTagging (ours) Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,388
@@
To address these issues, we present FinTagging, the first LLM-ready benchmark for full-scope, structure-aware XBRL tagging, which challenges models to perform end-to-end fact extraction and taxonomy alignment on corporate filings. As shown in Table 1, unlike prior work shah2023finer ; loukas2022finer ; sharma2023financial that frames tagging as flat multi-class classification and focuses solely on narrative text, FinTagging requires models to jointly extract financial facts and align them with the US-GAAP taxonomy across both unstructured text and structured tables. This is the first LLM-oriented formulation that scales to the full set of 10k+ taxonomy labels while avoiding the intractability of single-step classification, enabling more fine-grained evaluation of model performance. In collaboration with financial reporting experts, we decompose the task into two components: the first, structured numerical information extraction, assesses a model’s ability to identify key financial facts from both text and tables; the second, fine-grained concept linking, evaluates its ability to map each fact to the correct taxonomy concept among all semantically similar candidates. Unlike prior benchmarks limited to 1,000 frequent concepts, our setting covers the entire taxonomy, exposing the limitations of frequency-biased classification and demanding precise semantic reasoning bovee2005financial . We construct two evaluation sets: FinNI-eval  [^4] https://huggingface.co/datasets/TheFinAI/FinNI-eval for numerical fact extraction and FinCL-eval [^5] https://huggingface.co/datasets/TheFinAI/FinCL-eval for concept linking, both derived from real XBRL submissions and annotated with gold-standard mappings. We also introduce a unified evaluation framework that jointly measures extraction accuracy and concept alignment, providing a rigorous zero-shot assessment of LLMs’ capabilities for practical, table-aware XBRL tagging.
@@
We evaluate ten state-of-the-art LLMs under a zero-shot setting on three fronts: (1) end-to-end macro-F1 over the full FinTagging benchmark, (2) subtask-specific performance on FinNI and FinCL, and (3) ablation of our unified extraction-and-alignment evaluation framework. DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt achieve the highest macro-F1 scores, indicating that our benchmark design enables strong handling of both frequent and rare financial tags, an improvement over traditional PLM baselines evaluated under token-classification settings. In subtask analyses, these LLMs excel at numerical information extraction (FinNI) but continue to struggle with precise concept linking (FinCL), underscoring the remaining challenge of fine-grained semantic alignment. Crucially, our ablation study shows that, without the joint evaluation framework, even top LLMs produce invalid tagging outputs, highlighting the framework’s essential role in realistic assessment of XBRL tagging readiness.
@@
We conclude our main contributions as follows: (1) We introduce FinTagging, the first LLM-ready benchmark for full-scope, table-aware XBRL tagging with a two-stage pipeline—so that modern LLMs can be probed in true zero-shot setting. For each stage we release a new, professionally annotated evaluation set: FinNI-eval and FinCL-eval. (2) We conduct an extensive evaluation of state-of-the-art LLMs in zero-shot settings, systematically analyzing their performance across information extraction and semantic alignment subtasks, as well as overall tagging accuracy. (3) Our experimental results reveal a substantial performance gap between LLMs and the task requirements, particularly in fine-grained semantic alignment. This highlights the limitations of current LLMs in complex financial applications and underscores the need for continued advancements in task-specific adaptation.
2 Related Work
XBRL Tagging Benchmarks
@@ 2 Related Work
Previous benchmarks have framed XBRL tagging as a large-scale classification problem, primarily focusing on entity recognition. FiNER loukas2022finer introduced a dataset of 1.1 million sentences from SEC filings, annotated with 139 XBRL entity types, addressing challenges like context-sensitive numeric entities and financial domain-specific expressions. Building on this, Sharma et al. sharma2023financial proposed the Financial Numeric Extreme Labelling (FNXL) dataset, expanding the label space to approximately 2,800 XBRL tags and employing extreme classification techniques to handle the increased scale. While these works emphasize the complexity of financial taxonomies, their flat extreme classification approach overlooks the reasoning and alignment needed for realistic XBRL tagging, limiting their suitability for evaluating large language models.
XBRL Tagging Methods
@@
To improve XBRL tagging accuracy, prior work has incorporated structured knowledge and domain-specific modeling. Saini et al. saini2021galaxc proposed GalaXC, a graph neural network leveraging label hierarchies for improved classification. Ma et al. ma2022label enriched transformer models with taxonomy definitions to disambiguate semantically similar tags. Loukas et al. loukas2022finer introduced SECBERT and numeral-aware input transformations to enhance robustness. For extreme multi-label tagging, Sharma et al. sharma2023financial applied AttentionXML to focus on context-relevant segments. Wang wang2023standardizing aligned custom tags with standard taxonomies via semantic similarity using TF-IDF, Word2Vec, and FinBERT. More recently, Han et al. han2024xbrlagent developed XBRL-Agent, an LLM-based system for extracting insights from filings. These efforts highlight the shift toward taxonomy-aware and LLM-enhanced solutions for XBRL tagging.
Financial Evaluation Benchmarks
@@
Broader financial NLP benchmarks address information extraction, reasoning, and document understanding tasks. FiNER-ORD shah2023finer and FinRED sharma2022finred focus on entity recognition and relation extraction, respectively. BizBench koncel2023bizbench assesses the quantitative-reasoning ability of LLMs for both business and finance. Pixiu xie2023pixiu evaluates LLMs across classification, QA, and summarization, while FinQA chen2021finqa and ConvFinQA chen2022convfinqa focus on numerical reasoning over text and tables. TAT-QA zhu2021tat and DocVQA mathew2021docvqa target table-text joint understanding. These benchmarks highlight the importance of integrating structured and unstructured information, yet they overlook the taxonomy-driven fact alignment and do not support the structured output required for XBRL tagging.
3 FinTagging
3.1 Task Formulation
[…]

### v2

@@
Abstract.
Accurately understanding numbers from financial reports is fundamental to how markets, regulators, algorithms, and normal people read the economy and the world, yet even with XBRL (eXtensible Business Reporting Language) designed to tag every figure with standardized accounting concepts, mapping thousands of facts to over 10,000 U.S. GAAP concepts remains costly, inconsistent, and error-prone. Existing benchmarks define tagging as flat, single-step, extreme classification over small subsets of US-GAAP concepts, overlooking both the taxonomy’s hierarchical semantics and the structured nature of real tagging, where each fact must be represented as a contextualized multi-field output. These simplifications prevent fair evaluation of large language models (LLMs) under realistic reporting conditions. To address these gaps, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first comprehensive benchmark for structure-aware and full-scope XBRL tagging, designed to evaluate LLMs’ ability to extract and align financial facts through numerical reasoning and taxonomy alignment across text and tables. We define two subtasks: FinNI for numeric identification, which extracts numerical entities and their types from XBRL reports, and FinCL for concept linking, which maps each extracted entity to the corresponding concept in the full US-GAAP taxonomy. Together, these subtasks produce a structured representation of each financial fact. We evaluate diverse LLMs under zero-shot settings and analyze their performance across both subtasks and overall tagging accuracy. Results show that LLMs generalize well in numeric identification but struggle with fine-grained concept linking, revealing current limitations in structure-aware reasoning for accurate financial disclosure. All code and datasets are available on GitHub [^1] https://github.com/The-FinAI/FinTagging/ and Hugging Face [^2] https://huggingface.co/collections/TheFinAI/fintagging-68270132372c6608ac069bef.
XBRL tagging, Benchmark, Large language model, Information extraction, Information retrieval, Reranking
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Test collectionsccs: Information systems Information extractionccs: Information systems Clustering and classification
1. Introduction
@@
Automated tagging is essential in financial reporting, converting structured content such as tables and text into machine-readable data by linking each number to its meaning. Globally, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, which provide essential information for investors, regulators, and analysts. However, inconsistent terminology poses challenges for reliable interpretation. To address this, the eXtensible Business Reporting Language (XBRL) was introduced in 1999 as a global standard for machine-readable financial data (Richards et al., 2006). By applying structured tags, as shown in Figure 1, XBRL transforms fragmented disclosures into a consistent format, enabling seamless data exchange, cross-entity comparisons, and large-scale analysis. However, accurately tagging over 2,000 numerical facts per report to more than 10,000 standardized concepts remains labor-intensive and error-prone. In 2023 alone, over 6,500 tagging errors were identified across 33,000 SEC filings [^3] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
Figure 1. An example of Financial Tagging in Realistic, where the orange numbers mark facts to be tagged, black dashed arrows show the structured outputs after tagging, and blue dashed arrows denote the US-GAAP taxonomy referenced during tagging.
@@
Existing approaches to automated XBRL tagging rely on task-specific models and pretrained language models (PLMs) (Loukas et al., 2022; Sharma et al., 2023), but they require costly annotations, struggle to generalize to new filings, and perform poorly as taxonomies scale. In contrast, recent large language models (LLMs) show strong zero- and few-shot reasoning without task-specific supervision (Wei et al., 2022; Cao et al., 2025; Wang et al., 2025; Peng et al., 2025), yet their application to XBRL tagging remains underexplored due to two major limitations (Table 1). First, task framing is oversimplified: prior benchmarks like FiNER (Loukas et al., 2022) and FNXL (Sharma et al., 2023) treat tagging as extreme multi-class classification over flat label sets, offering little context for disambiguating fine-grained financial facts. These datasets typically cover only 1k+ concepts, leaving most of the 10k+ US-GAAP taxonomy untested. As shown in Table 8, SOTA LLMs such as DeepSeek-V3 (Liu et al., 2024) and GPT-4o (Hurst et al., 2024) achieve 0.0 precision, recall, and F1 under this setting. Second, structured data is ignored: existing datasets exclude tables, despite their central role in financial reporting. Many key facts appear in structured formats, as shown in recent financial QA benchmarks (Chen et al., 2021; Zhu et al., 2021; Mathew et al., 2021), making this omission a key gap between benchmarks and real-world tagging needs.
@@
Table 1. Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities. “Num.Reasoning” indicates the numerical reasoning. “Struct.IE” denotes the structured information extraction.
Benchmark Scenario Data Source Task Modality
#Entity
Label
#Taxonomy
Label
Num.
Reasoning?
Struct.
IE?
Concept
Linking?
FinQA (Chen et al., 2021) Decision making SEC 10-K QA text/table 0 0
ConvFinQA (Chen et al., 2022) Decision making SEC 10-K QA text/table 0 0
TAT-QA (Zhu et al., 2021) Financial analysis Chinese financial reports QA text/table 0 0
@@
DocVQA (Mathew et al., 2021) Enterprise automation Financial document images QA text/image 0 0
@@
FiNER-ORD (Shah et al., 2023) Financial tagging Chinese financial disclosures Classification text 49 0
@@
FinRED (Sharma et al., 2022) Knowledge graph construction English financial news Classification text 38 0
FiNER (Loukas et al., 2022) XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,388
FNXL (Sharma et al., 2023) XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,388
@@
FinTagging (ours) Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,388 / 17,388
@@
To address these issues, we present 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-ready benchmark for full-scope, structure-aware XBRL tagging, which challenges models to perform end-to-end fact extraction and taxonomy alignment on corporate filings. As shown in Table 1, unlike prior work (Shah et al., 2023; Loukas et al., 2022; Sharma et al., 2023) that frames tagging as flat multi-class classification and focuses solely on narrative text, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: requires models to jointly extract financial facts and align them with the US-GAAP taxonomy across both unstructured text and structured tables. This is the first LLM-oriented formulation that scales to the full set of 10k+ taxonomy labels while avoiding the intractability of single-step classification, enabling more fine-grained evaluation of model performance. In collaboration with financial reporting experts, we decompose the task into two components: the first, structured numerical information extraction, assesses a model’s ability to identify key financial facts from both text and tables; the second, fine-grained concept linking, evaluates its ability to map each fact to the correct taxonomy concept among all semantically similar candidates. Unlike prior benchmarks limited to 1,000 frequent concepts, our setting covers the entire taxonomy, exposing the limitations of frequency-biased classification and demanding precise semantic reasoning (Bovee et al., 2005). We construct two evaluation sets: FinNI-eval [^4] https://huggingface.co/datasets/TheFinAI/FinNI-eval-v1 for numerical fact extraction and FinCL-eval [^5] https://huggingface.co/datasets/TheFinAI/FinCL-eval-v1 for concept linking, both derived from real XBRL submissions and annotated with gold-standard mappings. We also introduce a unified evaluation framework that jointly measures extraction accuracy and concept alignment, providing a rigorous zero-shot assessment of LLMs’ capabilities for practical, table-aware XBRL tagging.
@@
We evaluate 13 state-of-the-art LLMs under a zero-shot setting on three fronts: (1) end-to-end macro-F1 over the full 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, (2) subtask-specific performance on FinNI and FinCL, and (3) ablation of our unified extraction-and-alignment evaluation framework. DeepSeek-V3 (Liu et al., 2024) and GPT-4o (Hurst et al., 2024) achieve the highest macro-F1 scores, indicating that our benchmark design enables strong handling of both frequent and rare financial tags, an improvement over traditional PLM baselines evaluated under token-classification settings. In subtask analyses, these LLMs excel at numerical information extraction (FinNI) but continue to struggle with precise concept linking (FinCL), underscoring the remaining challenge of fine-grained semantic alignment. Crucially, our ablation study shows that, without the joint evaluation framework, even top LLMs produce invalid tagging outputs, highlighting the framework’s essential role in realistic assessment of XBRL tagging readiness.
@@
We conclude our main contributions as follows: (1) We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-ready benchmark for full-scope, table-aware XBRL tagging with a two-stage pipeline, so that modern LLMs can be probed in a true zero-shot setting. For each stage, we release a new, professionally annotated evaluation set: FinNI-eval and FinCL-eval. (2) We conduct an extensive evaluation of state-of-the-art LLMs in zero-shot settings, systematically analyzing their performance across information extraction and semantic alignment subtasks, as well as overall tagging accuracy. (3) Our experimental results reveal a substantial performance gap between LLMs and the task requirements, particularly in fine-grained semantic alignment. This highlights the limitations of current LLMs in complex financial applications and underscores the need for continued advancements in task-specific adaptation.
2. Related Work
@@
XBRL Tagging Benchmarks: Previous benchmarks have framed XBRL tagging as a large-scale classification problem, primarily focusing on entity recognition. FiNER (Loukas et al., 2022) introduced a dataset of 1.1 million sentences from SEC filings, annotated with 139 XBRL entity types, addressing challenges like context-sensitive numeric entities and financial domain-specific expressions. Building on this, Sharma et al. (Sharma et al., 2023) proposed the Financial Numeric Extreme Labelling (FNXL) dataset, expanding the label space to approximately 2,800 XBRL tags and employing extreme classification techniques to handle the increased scale. While these works emphasize the complexity of financial taxonomies, their flat extreme classification approach overlooks the reasoning and alignment needed for realistic XBRL tagging, limiting their suitability for evaluating large language models.
@@
XBRL Tagging Methods: To improve XBRL tagging accuracy, prior work has incorporated structured knowledge and domain-specific modeling. Saini et al. (Saini et al., 2021) proposed GalaXC, a graph neural network leveraging label hierarchies for improved classification. Ma et al. (Ma et al., 2022) enriched transformer models with taxonomy definitions to disambiguate semantically similar tags. Loukas et al. (Loukas et al., 2022) introduced SECBERT and numeral-aware input transformations to enhance robustness. For extreme multi-label tagging, Sharma et al. (Sharma et al., 2023) applied AttentionXML to focus on context-relevant segments. Wang (Wang, 2023) aligned custom tags with standard taxonomies via semantic similarity using TF-IDF, Word2Vec, and FinBERT. More recently, Han et al. (Han et al., 2024) developed XBRL-Agent, an LLM-based system for extracting insights from filings. These efforts highlight the shift toward taxonomy-aware and LLM-enhanced solutions for XBRL tagging.
@@
Financial Evaluation Benchmarks: Broader financial NLP benchmarks address information extraction, reasoning, and document understanding tasks. FiNER-ORD (Shah et al., 2023) and FinRED (Sharma et al., 2022) focus on entity recognition and relation extraction, respectively. BizBench (Koncel-Kedziorski et al., 2023) assesses the quantitative-reasoning ability of LLMs for both business and finance. Pixiu (Xie et al., 2023) evaluates LLMs across classification, QA, and summarization, while FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) focus on numerical reasoning over text and tables. TAT-QA (Zhu et al., 2021) and DocVQA (Mathew et al., 2021) target table-text joint understanding. These benchmarks highlight the importance of integrating structured and unstructured information, yet they overlook the taxonomy-driven fact alignment and do not support the structured output required for XBRL tagging.
3. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
3.1. Task Formulation
Formally, given a financial document $D$ consisting of multiple textual contents and structured tables, a set of predefined value data types $L$ specified by the XBRL specification, and a taxonomy database $\mathcal{T}$ containing a set of financial semantic concepts, the XBRL tagging task is to identify all relevant numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. Here, Tag denotes the linked concept in $\mathcal{T}$ (e.g. “us-gaap:CashAndDueFrom Banks”, etc.), Fact represents the extracted numerical value as it appears in context (e.g., “2.5”, “10,000”, “two”, etc.), and Type specifies the corresponding data type in $L$ (including monetaryItemType, percentItemType, sharesItemType, perShareItem, and integerItemType). Each numerical fact is assigned a single, most specific tag from the taxonomy, without overlapping. Specifically, monetaryItemType refers to a financial amount, such as revenue or expenses, typically reported in currency units (e.g., USD). percentItemType represents a rate or ratio, expressed as a decimal (e.g., 25%). sharesItemType indicates the number of shares, such as stocks held or issued. perShareItemType captures values reported on a per-share basis, like earnings per share. Lastly, integerItemType is used for whole number counts, such as the number of employees or transactions.